a car travels from one town to another at a speed of 18mph. if it had gone 18 mph faster, it could of made the trip in a half hour less time. how far apart are the towns?
Answer by mananth(16949)

let the distance between towns be x miles
initial speed = 18 mph
time taken = x/18
..
if speed is increased by 18
then speed will be 36 mph
distance =x
time = x/36
x/8- x/36=1/2
LCD= 72
9x-2x / 72=1/2
7x/72=1/2
7x= 72/2
x=72/2*7
x=5.14 miles